# MACTAS: Self-Attention-Based Module for Inter-Agent Communication in Multi-Agent Reinforcement Learning

### v1

@@ C.3 MMM2
C.5 3s5z_vs_3s6z
Figure 14: 3s5z_vs_3s6z replay.
MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot. The stalkers are fast, range-attacking units, but with weak defense. On the other hand, zealots have good defense, but need to get close to the opponent unit. The tactic in this scenario is interesting. Using one stalker as a decoy, MACTAS first defeats the enemy stalkers and then proceeds to eliminate the rest of the enemy force. One agent, a stalker, sacrifices himself and focuses a few enemy zealots on himself. In Figure 14, we can see 5 enemy zealots following the stalker marked as “Local Player”. The remaining MACTAS’ 2 stalkers and 5 zealots attack 3 enemy stalkers and 1 zealot with an advantage. This move results in the elimination of enemy stalkers and the zealot. In the end, MACTAS’ units eliminate zealots lured by the sacrificed stalker.
@@ Appendix D Sensitivity analysis
Figure 15: MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
Figure 16: MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
Figure 17: MACTAS; sensitivity of the results to architecture size with the VDN mixer. Due to the technical issues, the version with 3 Transformer layers and a width of the Transformer feedforward layer equal to 512 was evaluated on 4 seeds of the random generator (1 to 4).
Figure 18: MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
Figure 19: MACTAS; sensitivity of the top 2 exploration method to the temperature coefficient.
We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer in five environments: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2 and 6h_vs_8z. We experimented with the dropout of the Transformer encoder on all six maps. We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on three maps: 27m_vs_30m, MMM2 and 3s5z_vs_3s6z.
@@
Figure 19 shows the impact of the softmax temperature parameter on the proposed exploration scheme. We observe that this method is sensitive to this parameter. For the 27m_vs_30m and MMM2 maps, the low temperatures work better, with $0.25$ achieving similar results to the $\epsilon$-greedy scheme, which is equivalent to choosing 0 as the temperature. In these cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration. We can observe an interesting situation on the 3s5z_vs_3s6z map. The top 2 exploration scheme outperforms the $\epsilon$-greedy exploration for the $0.33$ temperature. However, a temperature equal to $4$ gives a similar result to the $\epsilon$-greedy exploration. The in-between temperatures perform worse than both of these values, showing that there is no monotonic relationship between the temperature and the performance.
@@ Appendix E Software
Table 3 contains a list of key programming libraries used during our experiments. The complete list of the programming libraries is available in the MACTAS’ official code repository in file requirements.txt.
@@
Table 3: Essential libraries in programming environment.
@@ Appendix F Hardware
Table 4: Specification of the computational clusters used in the experiments.
Here we present the hardware used in our experiments. We used NVIDIA Ampere GPU architecture on three computational clusters. The specific setup is presented in Table 4.

### v2

@@ C.3. MMM2
C.5. 3s5z_vs_3s6z
Figure 15. 3s5z_vs_3s6z replay.
MACTAS controls 3 stalkers and 5 zealots, while the enemy has one more zealot. The stalkers are fast, range-attacking units, but with weak defense. On the other hand, zealots have a good defense, but they need to get close to the opponent’s unit. The tactic in this scenario is interesting. Using one stalker as a decoy, MACTAS first defeats the enemy stalkers and then proceeds to eliminate the rest of the enemy force. One agent, a stalker, sacrifices himself and focuses a few enemy zealots on himself. In Figure 15, we can see 5 enemy zealots following the stalker marked as “Local Player”. The remaining MACTAS’ 2 stalkers and 5 zealots attack 3 enemy stalkers and 1 zealot with an advantage. This move results in the elimination of enemy stalkers and the zealot. In the end, MACTAS’ units eliminate zealots lured by the sacrificed stalker.
@@ Appendix D Sensitivity analysis
Figure 16. MACTAS; sensitivity of the results to architecture size with the QMIX mixer.
Figure 17. MACTAS; sensitivity of the results to architecture size with the QPLEX mixer.
Figure 18. MACTAS; sensitivity of the results to architecture size with the VDN mixer.
Figure 19. MACTAS; sensitivity of the results to the dropout parameter in the Transformer block. The default value of this parameter is $10\%$.
Figure 20. MACTAS; sensitivity of the top 2 exploration methods to the temperature coefficient.
We evaluated the choice of the number of Transformer encoder layers and the width of the Transformer feedforward layer for MACTAS with QMIX, QPLEX, and VDN as a mixer on six SMAC(v1) maps: 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, 6h_vs_8z, and 3s5z_vs_3s6z. We experimented with the dropout of the Transformer encoder on all six maps. We also assessed the sensitivity of our new top $k$ exploration technique to the softmax temperature in the case of MACTAS+QMIX and $k=2$ on all six maps.
@@
Figure 20 shows the impact of the softmax temperature parameter on the proposed exploration scheme. We observe that this method is sensitive to this parameter. For the 2c_vs_64zg, 5m_vs_6m, 27m_vs_30m, MMM2, and 6h_vs_8z maps, the low temperatures ($0.25$, $0.33$) work best with better results for the $0.25$ temperature. We can observe an interesting situation on the 3s5z_vs_3s6z map. The best results are obtained for the $0.33$ temperature, with the temperature $4$ being the second-best, and the temperature $0.25$ performing quite poorly, which suggests that, in general, the relationship between the temperature value and performance may not be monotonic. For the most challenging maps, which require 7.5 million training steps, the top 2 exploration schemes outperform the $\epsilon$-greedy exploration, which is equivalent to choosing 0 as the temperature. In other cases, the new exploration scheme does not outperform the $\epsilon$-greedy exploration; however, except for the case of the 5m_vs_6m map, where training is slowed down, the method remains indifferent.
@@ Appendix E Software
Table 3 contains a list of key programming libraries used during our experiments. The complete list of programming libraries is available in the MACTAS’ official code repository in the file requirements.txt.
@@
Table 3. Essential libraries in programming environment.
@@ Appendix F Hardware
Table 4. Specification of the computational clusters used in the experiments.
Here we present the hardware used in our experiments. We utilized the NVIDIA Ampere GPU architecture across three computational clusters. The specific setup is presented in Table 4.